# ExLLM: Experience-Enhanced LLM Optimization for Molecular Design and Beyond

### v4

@@ 6 Conclusion
We presented ExLLM, an LLM-as-optimizer framework for large discrete optimization, instantiated for multi-objective molecular design. The method couples three simple components: a single, evolving experience that distills non-redundant cues at low cost; a $k$-offspring sampling scheme that widens exploration per query; and a feedback adapter that normalizes objectives for selection while formatting constraints and expert hints for iteration. Under a fixed evaluation budget, ExLLM attains state-of-the-art results on PMO (total score 19.165, ranking first on 17/23 tasks, +7.3% over prior SOTA) and generalizes beyond chemistry, setting new records on circle packing and stellarator design while delivering strong performance across additional domains. The probabilistic experience injection improves efficiency without over-constraining exploration. Looking forward, we plan to (i) adapt $k$ and the experience-injection rate online, (ii) incorporate oracle uncertainty and 3D/physics-informed feedback, and (iii) broaden evaluation to additional scientific design tasks. We release templates for rapid transfer, lowering the barrier to applying ExLLM as a general optimizer for large discrete spaces.
@@ Acknowledgements
Author Contributions. Nian Ran led the algorithm design, implementation, and experiments. Yue Wang supervised the research, provided computational resources, and contributed to paper structure and revision. Xiaoyuan Zhang, Zhongzheng Li, Qingsong Ran, and Wenhao Li contributed to experiments and application studies. Richard Allmendinger provided research guidance and feedback on paper structure and presentation.

### v5

@@ 5 Conclusion
We presented ExLLM, an LLM-as-optimizer framework for large discrete optimization, instantiated for multi-objective molecular design. The method couples three simple components: a single, evolving experience that distills non-redundant cues at low cost; a $k$-offspring sampling scheme that widens exploration per query; and a feedback adapter that normalizes objectives for selection while formatting constraints and expert hints for iteration. Under a fixed evaluation budget, ExLLM attains state-of-the-art results on PMO (total score 19.165, ranking first on 17/23 tasks, +7.3% over prior SOTA) and generalizes beyond chemistry, setting new records on circle packing and stellarator design while delivering strong performance across additional domains including offshore jacket optimization, MOCPOP benchmark, peptide design and GCU kernel optimization. The probabilistic experience injection improves efficiency without over-constraining exploration. Looking forward, we plan to (i) adapt $k$ and the experience-injection rate online, (ii) incorporate oracle uncertainty and 3D/physics-informed feedback, and (iii) broaden evaluation to additional scientific design tasks. We release templates for rapid transfer, lowering the barrier to applying ExLLM as a general optimizer for large discrete spaces.
@@ Acknowledgements
Author Contributions: Nian Ran led the algorithm design, implementation, and experiments. Yue Wang supervised the research, provided computational resources, and contributed to paper structure and revision. Xiaoyuan Zhang, Zhongzheng Li, Qingsong Ran, and Wenhao Li contributed to experiments and application studies. Richard Allmendinger provided research guidance and feedback on paper structure and presentation.